OpenStudy (heywhatsupheyo):

Help me please. I understand most of the concept but im still having difficulty. x^2 - 9 =0 we're solving quadratics by completing the square using the square root property

OpenStudy (will.h):

Okay. What is the square root of x^2?

OpenStudy (will.h):

And what is the square root of 9

OpenStudy (heywhatsupheyo):

would the square root of x^2 be x and the square root of 9 be 3?

OpenStudy (will.h):

Indeed Now form parentheses to gather the data that we have so far (x 3) (x 3) = 0 What would be the sign in between those terms? Positive or negative for the 1st parentheses and positive or negative for the second?

OpenStudy (will.h):

Hint In the original expression X^2 - 9 We have a negative sign in between. That would lead to eliminate the term in the middle which would require both positive and negative signs.. Any ideas?

OpenStudy (heywhatsupheyo):

would it be + for the first set of parentheses and - for the second set? such as (x+3)(x-3)

OpenStudy (will.h):

It would be.. For one of them (x -3) (x+ 3) = 0 See that if we get to simplify this We would get the following, after applying destrubtive property x * x = x^2 X * 3 = 3x -3 * x = -3x -3 * 3 = -9 Which means x^2 + 3x -3x -9 = 0 Notice that the terms in the middle 3x and -3x can be eliminated which would leave us with x^2 -9 = 0 See that we got the same original expression which means our answer is right. That's one way to check your work

OpenStudy (heywhatsupheyo):

oh okay.